# every label/.append style = font =

### v1

@@ Abstract
We propose a superconducting diode device comprising a central superconducting film flanked by two wires carrying an applied DC bias, suitably chosen so as to generate different asymmetric field profiles. Through numerical simulations of the coupled Ginzburg-Landau and heat-diffusion equations, we show that this design is capable of efficiently breaking the reciprocity of the critical current in the central superconductor, thus promoting the diode effect in response to an applied AC current. By adjusting the DC bias in the wires, we find the optimum inhomogeneous field profile that facilitates the entrance of vortices and antivortices in a given polarity of the applied AC current and impede their entrance in the other polarity. This way, the system behaves as an ideal superconducting half-wave rectifier, with diode efficiencies surpassing 70%. Furthermore, we detail the behavior and diode efficiency of the system under different experimental conditions, such as the substrate heat transfer coefficient and the sweep rate of the external current.
@@ I Introduction
At the core of the superconducting diode lies a fundamental characteristic, the absence of symmetry in current flow, expressed as $I_{c}^{+}\neq I_{c}^{-}$, where $I_{c}^{\pm}$ is the critical current leading to the resistive state at a given current direction. This concept originated decades ago as a mean of manipulating the density and motion of magnetic flux quanta, known as vortices, within superconductors [3, 4]. The primary goal was to disrupt the inversion symmetry of the pinning potential acting on vortices. This would enable an alternating current to induce vortex motion solely along the potential’s easy direction, resulting in a resistive flux-flow state during one half cycle and purely superconducting current flow during the other one. Subsequently, experimental realization of these flux quanta diodes swiftly followed, with diverse research groups employing various nanolithographic techniques to selectively break spatial inversion symmetry for vortex motion, such as regular arrays of asymmetric dots and antidots [5, 6, 7], asymmetric distributions of symmetric dots and antidots [8, 9, 10], regular arrays of nanosized magnetic dipoles [11, 12], and asymmetric weak-pinning channels [13].
@@
Intrinsic superconducting diode effects (SDE) rely on non-reciprocal pair-breaking mechanisms inherent to the material [39, 40, 41]. Although these mechanisms are of fundamental importance, they lack the high degree of controllability of vortex-based diode effects (VDE), where the non-reciprocal critical currents can be carefully designed by e.g. nanostructuring. Furthermore, in typical experimental situations, the superconducting sample enters the resistive state at currents considerably lower than the theoretical depairing current, either as a result of vortex motion or the nucleation of supercritical hot spots. This makes it challenging to discern experimentally intrinsic and non-intrinsic effects [42].
@@
During the resistive state of a VDE, the material is still superconducting and dissipation is induced by vortex motion (flux-flow). Since the flux-flow resistance is necessarily lower than the normal state resistance, the measured DC signals in VDE are relatively small. This is a clear disadvantage when compared to the strict superconducting diode, where the dissipationless superconducting state for current applied in one direction is expected to transition directly to the fully normal state when the current is reversed, leading to a considerably larger DC signal. Remarkably, Hou et al. [16] recently reported on the observation of highly efficient SDE on conventional superconducting films where the AC switching between superconducting and normal states is induced by non-reciprocal penetration of vortices due to asymmetrical conditions at the sample edges, that is, by an extrinsic, vortex-based mechanism. However, the reason why the flux-flow state was suppressed in favor of a sudden jump to the full normal state, thus making this VDE into a SDE, is unclear. A possible explanation can be inferred from Ref. [43], where the authors investigated the diode effect in a superconducting stripe with a conformal array of nanoholes of broken inversion symmetry. This work suggests that hot posts induced by fast vortex motion can drive the superconductor directly into the normal state thus bridging the gap between VDE and SDE.
@@
These recent advances emphasize the necessity of better understanding the role of heating effects in superconducting diodes. Additionally, the suggestion that the VDE can be made into a fully SDE redirects the research on vortex diode systems towards the optimization of the diode efficiency, defined as $\epsilon=|I_{c}^{-}-I_{c}^{+}|/(I_{c}^{-}+I_{c}^{+})$. Tackling these challenges is a key point for obtaining more efficient and reliable superconducting diodes.
@@
In the present work, we investigate the interplay between the superconducting and the vortex diode effects (SDE and VDE) in thin superconducting films exposed to external asymmetric fields. By integrating the coupled time-dependent Ginzburg-Landau and heat diffusion equations, we explore simple inhomogeneous field landscapes, from asymmetric fully positive to antisymmetric (half-positive, half-negative) flux profiles, induced either by ferromagnetic films or DC currents in nearby wires as illustrated in Fig. 1. Our findings reveal that both profile types promote VDE, with the antisymmetric one exhibiting the highest diode efficiencies, up to 70%. Furthermore, at large enough AC amplitudes, VDE is replaced by SDE as a result of vortex-antivortex collisions and subsequent hotspot proliferation. We further demonstrate that reducing the substrate’s heat transfer coefficient accelerates the VDE-to-SDE transition, thereby enhancing the superconducting diode efficiency. These insights illuminate recent discoveries of highly efficient SDE and emphasize the role of vortex-antivortex collisions and the importance of optimizing substrate designs and cooling environments for future superconducting diode platforms.
@@ II Theoretical Formalism
Eqs. (1)-(3) are solved numerically only for the central superconductor, supplemented by periodic boundary conditions along $x$, with period $L_{x}=400\xi(0)$. At the edges $y=\pm L_{y}/2$, $\psi$ and $T$ satisfy the boundary conditions $\hat{\textbf{y}}\cdot(\mbox{\boldmath$\nabla$}-i\textbf{A})\psi=0$ and $\hat{\textbf{y}}\cdot\mbox{\boldmath$\nabla$}T=0$ [46, 47]. Additionally, the local field $\textbf{h}=\bm{\nabla}\times\textbf{A}$ at the sample edges satisfies $h_{z}(x,\pm L_{y}/2)=H_{nh}(\pm L_{y}/2)\pm L_{y}J(t)/2\kappa^{2}$, which formally accounts for the applied current $J(t)$ in the sample [48].
@@ III Results and Discussion
To analyze the conditions for the occurrence of the superconducting diode effect or the vortex diode effect in our system, we have varied both the amplitude of the applied AC current and the values of $J_{1}$ and $J_{2}$ carried by each side stripe. We limit $J_{1}$ and $J_{2}$ to be lower than $0.52~{}\sigma\hbar/2e\xi(0)t_{GL}$. With the thermal bath temperature fixed at $T_{0}=0.96T_{c}$, such currents would destroy superconductivity in the central stripe. Therefore, the side stripes must be composed of a superconducting material with larger critical temperature and depairing current density. In what follows, voltages are expressed in units of $V_{0}=\hbar/2et_{GL}$ and current densities in units of $j_{0}=\sigma V_{0}/\xi(0)$. Further, we fix $\kappa=5$ and, unless stated otherwise, the period of the AC current is $\tau=\tau_{0}=10^{5}t_{GL}$.
@@ III.2 Antisymmetric flux profiles ($J_{1}=J_{2}$)
the voltage output as a function of time for three different values of $J_{a}$ and $J_{1}=J_{2}=0.26$. As shown in panel $(a)$, for the positive branch, shortly after $t=0$, when the magnitude of the AC current is sufficient to put in motion the vortices induced by the inhomogeneous field, the voltage becomes finite and gradually increases with current. In this case, though, the dissipation process is never large enough to totally suppress the superconducting state and we get a behaviour that resembles the positive branch of panel $(a)$ in Fig. 3. On the other hand, the negative current branch is significantly different from the previous studied cases. As can be seen, after a small voltage signal corresponding to the vortex induced by the inhomogeneous field leaving the sample, the voltage becomes zero for the remaining half period of negative current. In other words, the system behaves as a perfect half-wave rectifier.
@@ III.4 Diode efficiency
Let us now investigate the efficiency of our superconducting diode device, which is defined as $\epsilon_{SDE}=\left(I_{c}^{-}-I_{c}^{+}\right)/\left(I_{c}^{-}+I_{c}^{+}\right)$, where $I_{c}^{\pm}$ is the critical current for the complete destruction of the superconducting state for a given current polarity. To do this, we replace the sine wave AC current by a linear current cycle, with total sweep time $\tau=10\tau_{0}=10^{5}t_{GL}$, thereby mimicking the usual experimental procedure of measuring the voltage in current-sweep mode. We keep $J_{1}=J_{2}=0.26$ fixed, once these parameters make possible the occurrence of the ideal superconducting diode scenario. Fig. 9 shows the voltage signal as a function of time for these parameters. $I_{c}^{\pm}$ is highlighted in this figure by the two blue arrows and their values give us $\epsilon_{SDE}=0.124$.
From Fig. 9 we can define another useful critical current, namely the current which marks the onset of vortex motion, $I_{cv}$, which are marked by the orange arrows. If we now define $\epsilon_{VDE}=\left(I_{cv}^{-}-I_{cv}^{+}\right)/\left(I_{cv}^{-}+I_{cv}^{+}\right)$ as the efficiency of the diode effect caused by the finite voltage associated with the flux-flow, we have $\epsilon_{VDE}=0.619$ for this set of parameters. The much larger value of $\epsilon_{VDE}$ compared to $\epsilon_{SDE}$ can be promptly understood if we notice that, for positive current polarity, the profile of the self-field produced by the AC current matches the profile of the external inhomogeneous field. As a result, a large amount of vortices is present in the superconductor, meaning that a small current amplitude is sufficient to start the flux-flow. In contrast, for negative current polarity, the profiles of the inhomogeneous field and the AC current self-field are opposites. This means that there are no vortices in the system, unless the current amplitude is large enough to overcome the Bean-Livingston barrier. This is depicted at the second orange arrow in Fig. 9 which occurs at a large current amplitude, very close to $I_{c}^{-}$. This discrepancy is the origin of the large value of $\epsilon_{VDE}$.
@@ III.5 Optimizing diode efficiencies: the role of heat removal and sweep rate
So far, when solving the heat diffusion equation, we have set $\eta=2\times 10^{-4}$, which mimics a substrate of moderate heat removal efficiency. Once we have seen that the destruction of the superconducting state is governed by the emergence and expansion of hot spots through the system, it is natural to assume that the efficiency of our diode system depends on the heat removal coefficient $\eta$. Fig. 11 shows both $\epsilon_{SDE}$ and $\epsilon_{VDE}$ as a function of $\eta$, for values raging from $2\times 10^{-5}$ (weak heat removal) to $2\times 10^{-3}$ (strong heat removal).
@@
Figure 11: Efficiencies of the superconducting, $\epsilon_{\text{SDE}}$ (blue circles), and vortex, $\epsilon_{\text{VDE}}$ (red circles), diode effects as a function of the heat transfer coefficient $\eta$. Here, we set $\tau=10^{6}t_{GL}$.
@@
The vortex diode efficiency presents a weak dependence on $\eta$, which is a result of the low heat dissipation produced by vortices when driven by currents just above $I_{cv}^{+}$ (see the low flux-flow resistance regime shown in Fig. 12). However, as the current increases above $I_{cv}^{+}$, the rate of v-av collisions at the equatorial line of the superconducting film increases, leading to the formation of hot spots as discussed in the previous subsection. For high $\eta$, weak hot spots are quickly removed by the substrate so that a current much larger than $I_{cv}^{+}$ is necessary to drive the system into the normal state. In contrast, for small $\eta$, even weak hot spots, created in the early stages of the flux-flow regime, can grow and drive the film into the normal state. This can considerably reduce the value of $I_{c}^{+}$, bringing it closer to $I_{cv}^{+}$. On the other hand, during the negative current branch, the reduction of $I_{c}^{-}$ induced by weak heat removal is much less pronounced because $I_{c}^{-}$ is already very close to $I_{cv}^{-}$, even in the strong heat removal regime. Combined, these results lead to a pronounced enhancement of $\epsilon_{SDE}$ from 0.067 to 0.317 as $\eta$ is reduced from $2\times 10^{-3}$ to $2\times 10^{-5}$. Therefore, we can conclude that an ideal superconducting diode system should have the weakest possible heat removal scenario.
@@
It is worth mentioning that our 2D simulations are limited to the assumption that the thermal healing length is smaller than the combined thicknesses of film and substrate. In actual experiments, this condition is not generally fulfilled and typical values of $\eta$ can be much smaller than those used in our simulations [46]. In those cases, hot spots can nucleate as soon as flux-flow commences, thus aligning $I_{c}^{+}$ to $I_{cv}^{+}$ and thereby $\epsilon_{SDE}$ to $\epsilon_{VDE}$. Bearing this in mind, we speculate that the sharp superconductor-to-normal transitions observed in Ref. [16] are probably triggered by hot spots induced by the fast penetration of vortices or vortex-antivortex annihilation events in a scenario of very small $\eta$.
@@
Another important factor that influences the efficiency of the diode device is the total sweep time $\tau$, which determines the sweep rate along the current cycle. Fig. 12 shows both $\epsilon_{SDE}$ and $\epsilon_{VDE}$ as a function of $\tau$, with $\eta=2\times 10^{-5}$ fixed. As can be seen, both efficiencies present similar behavior, rapidly increasing for small values of $\tau$ and then approaching a constant value for large values. By increasing $\tau$, we observe a decrease in both $I_{c}^{\pm}$ and $I_{cv}^{\pm}$. The large efficiency occurs because $I_{c}^{+}$ ($I_{cv}^{+}$) decreases in a larger rate than $I_{c}^{-}$ ($I_{cv}^{-}$). In this case, a faster increase rate of the current, i.e., a lower value of $\tau$, leads to larger $I_{c}$ and $I_{cv}$, once there is not enough time for the system to reach the stable state for each region of current values.
@@
Figure 12: Efficiencies of the superconducting, $\epsilon_{\text{SDE}}$ (blue circles), and vortex, $\epsilon_{\text{VDE}}$ (red circles), diode effects as a function of total sweep time $\tau$. Here, $\eta=2\times 10^{-5}$.
@@
III.6 Optimizing diode efficiencies: the role of the inhomogeneous field amplitude

### v2

# Harnessing the superconducting diode effect through inhomogeneous magnetic fields

@@ Abstract
We propose a superconducting diode device comprising a central superconducting film flanked by two wires carrying an applied DC bias, suitably chosen so as to generate different asymmetric field profiles. Through numerical simulations of the coupled Ginzburg-Landau and heat-diffusion equations, we show that this design is capable of efficiently breaking the reciprocity of the critical current in the central superconductor, thus promoting the diode effect in response to an applied AC current. By adjusting the DC bias in the wires, we find the optimum inhomogeneous field profile that facilitates the entrance of vortices and antivortices in a given polarity of the applied AC current and impede their entrance in the other polarity. This way, the system behaves as a superconducting half-wave rectifier, with diode efficiencies surpassing 70%. Furthermore, we detail the behavior and diode efficiency of the system under different experimental conditions, such as the substrate heat transfer coefficient and the sweep rate of the external current.
@@ I Introduction
At the core of the superconducting diode lies a fundamental characteristic, the absence of symmetry in current flow, expressed as $I_{c}^{+}\neq I_{c}^{-}$, where $I_{c}^{\pm}$ is the critical current leading to the resistive state at a given current direction. However, the superconducting diode concept originated decades ago as a mean of manipulating the density and motion of magnetic flux quanta, known as vortices, within superconductors [3, 4]. The primary goal was to disrupt the inversion symmetry of the pinning potential acting on vortices. This would enable an alternating current to induce vortex motion solely along the potential’s easy direction, resulting in a resistive flux-flow state during one half cycle and purely superconducting current flow during the other one. Subsequently, experimental realization of these flux quanta diodes swiftly followed, with diverse research groups employing various nanolithographic techniques to selectively break spatial inversion symmetry for vortex motion, such as regular arrays of asymmetric dots and antidots [5, 6, 7], asymmetric distributions of symmetric dots and antidots [8, 9, 10], regular arrays of nanosized magnetic dipoles [11, 12], and asymmetric weak-pinning channels [13].
@@
Intrinsic superconducting diodes rely on non-reciprocal pair-breaking mechanisms inherent to the material [39, 40, 41]. Although these mechanisms are of fundamental importance, they lack the high degree of controllability of vortex-based red superconducting diodes, where the non-reciprocal critical currents can be carefully designed by e.g. nanostructuring. Furthermore, in typical experimental situations, the superconducting sample enters the resistive state at currents considerably lower than the theoretical depairing current, either as a result of vortex motion or the nucleation of supercritical hot spots. This makes it challenging to discern experimentally intrinsic and non-intrinsic effects [42].
@@
During the resistive state of a vortex-based diode, the material is still superconducting and dissipation is induced by vortex motion (flux-flow). Since the flux-flow resistance is necessarily lower than the normal state resistance, the measured DC signals in this case are relatively small. This is a clear disadvantage when compared to diodes where the dissipationless superconducting state for current applied in one direction transitions directly to the fully normal state when the current is reversed, leading to a considerably larger DC signal. We shall refer to these two scenarios as flux-flow dominated (FFD) and normal-state dominated (NSD) diode effects, respectively. Remarkably, Hou et al. [16] recently reported on the observation of highly efficient diode effect on conventional superconducting films where the AC switching between superconducting and normal states is induced by non-reciprocal penetration of vortices due to asymmetrical conditions at the sample edges, that is, by an extrinsic, vortex-based mechanism. However, the reason why the flux-flow state was suppressed in favor of a sudden jump to the full normal state, thus making this FFD diode into a NSD diode, is unclear. A possible explanation can be inferred from Ref. [43], where the authors investigated the diode effect in a superconducting stripe with a conformal array of nanoholes of broken inversion symmetry. This work suggests that hot posts spots induced by fast vortex motion can drive the superconductor directly into the normal state thus bridging the gap between FFD and NSD scenarios.
@@
These recent advances emphasize the necessity of better understanding the role of heating effects in superconducting diodes. Additionally, the suggestion that a FFD diode can be made into a NSD diode redirects the research on vortex diode systems towards the optimization of the diode efficiency, defined as $\epsilon=|I_{c}^{-}-I_{c}^{+}|/(I_{c}^{-}+I_{c}^{+})$. Tackling these challenges is a key point for obtaining more efficient and reliable superconducting diodes.
@@
In the present work, we investigate the interplay between the flux-flow dominated (FFD) and normal-state dominated (NSD) operation modes of the superconducting diode effect in thin superconducting films exposed to external asymmetric fields. By integrating the coupled time-dependent Ginzburg-Landau and heat diffusion equations, we explore simple inhomogeneous field landscapes, from asymmetric fully positive to antisymmetric (half-positive, half-negative) flux profiles, induced either by ferromagnetic films or DC currents in nearby wires as illustrated in Fig. 1. Our findings reveal that both profile types promote the FFD mode, with the antisymmetric one exhibiting the highest diode efficiencies, up to 70%. Furthermore, at large enough AC amplitudes, FFD is replaced by NSD as a result of vortex-antivortex collisions and subsequent hotspot proliferation. We further demonstrate that reducing the substrate’s heat transfer coefficient accelerates the FFD-to-NSD transition, thereby enhancing the power output of the diode. These insights illuminate recent discoveries of highly efficient superconducting diodes and emphasize the role of vortex-antivortex collisions and the importance of optimizing substrate designs and cooling environments for future superconducting diode platforms.
@@ II Theoretical Formalism
Eqs. (1)-(3) are solved numerically only for the central superconductor, supplemented by periodic boundary conditions along $x$, with period $L_{x}=400\xi(0)$. In the Supplementary Material, we show that our choice of $L_{x}$ does not affect the physical behavior of the system. At the edges $y=\pm L_{y}/2$, $\psi$ and $T$ satisfy the boundary conditions $\hat{\textbf{y}}\cdot(\mbox{\boldmath$\nabla$}-i\textbf{A})\psi=0$ and $\hat{\textbf{y}}\cdot\mbox{\boldmath$\nabla$}T=0$ [46, 47]. Additionally, the local field $\textbf{h}=\bm{\nabla}\times\textbf{A}$ at the sample edges satisfies $h_{z}(x,\pm L_{y}/2)=H_{nh}(\pm L_{y}/2)\pm L_{y}J(t)/2\kappa^{2}$, which formally accounts for the applied current $J(t)$ in the sample [48].
@@ III Results and Discussion
To analyze the conditions for the occurrence of the superconducting diode effect or the vortex diode effect in our system, we have varied both the amplitude of the applied AC current and the values of $J_{1}$ and $J_{2}$ carried by each side stripe. We limit $J_{1}$ and $J_{2}$ to be lower than $0.52~{}\sigma\hbar/2e\xi(0)t_{GL}$. With the thermal bath temperature fixed at $T_{0}=0.96T_{c}$ [49], such currents would destroy superconductivity in the central stripe. Therefore, the side stripes must be composed of a superconducting material with larger critical temperature and depairing current density. In what follows, voltages are expressed in units of $V_{0}=\hbar/2et_{GL}$ and current densities in units of $j_{0}=\sigma V_{0}/\xi(0)$. Further, we fix $\kappa=5$ and, unless stated otherwise, the period of the AC current is $\tau=\tau_{0}=10^{5}t_{GL}$.
@@ III.2 Antisymmetric flux profiles ($J_{1}=J_{2}$)
the voltage output as a function of time for three different values of $J_{a}$ and $J_{1}=J_{2}=0.26$. As shown in panel $(a)$, for the positive branch, shortly after $t=0$, when the magnitude of the AC current is sufficient to put in motion the vortices induced by the inhomogeneous field, the voltage becomes finite and gradually increases with current. In this case, though, the dissipation process is never large enough to totally suppress the superconducting state and we get a behaviour that resembles the positive branch of panel $(a)$ in Fig. 3. On the other hand, the negative current branch is significantly different from the previous studied cases. As can be seen, after a small voltage signal corresponding to the vortex induced by the inhomogeneous field leaving the sample, the voltage becomes zero for the remaining half period of negative current. In other words, the system behaves as a half-wave rectifier.
@@ III.4 Diode efficiency
Let us now investigate the efficiency of our superconducting diode device when operating in the NSD mode, defined as $\epsilon_{NSD}=\left(I_{c}^{-}-I_{c}^{+}\right)/\left(I_{c}^{-}+I_{c}^{+}\right)$, where $I_{c}^{\pm}$ is the critical current for the complete destruction of the superconducting state for a given current polarity. To do this, we replace the sine wave AC current by a linear current cycle, with total sweep time $\tau=10\tau_{0}=10^{5}t_{GL}$, thereby mimicking the usual experimental procedure of measuring the voltage in current-sweep mode. We keep $J_{1}=J_{2}=0.26$ fixed, once these parameters make possible the occurrence of the superconducting diode scenario with largest efficiency. Fig. 9 shows the voltage signal as a function of time for these parameters. $I_{c}^{\pm}$ is highlighted in this figure by the two blue arrows and their values give us $\epsilon_{NSD}=0.124$.
From Fig. 9 we can define another useful critical current, namely the current which marks the onset of vortex motion, $I_{cv}$, which are marked by the orange arrows. If we now define $\epsilon_{FFD}=\left(I_{cv}^{-}-I_{cv}^{+}\right)/\left(I_{cv}^{-}+I_{cv}^{+}\right)$ as the efficiency of the diode effect caused by the finite voltage associated with the flux-flow, we have $\epsilon_{FFD}=0.619$ for this set of parameters. The much larger value of $\epsilon_{FFD}$ compared to $\epsilon_{NSD}$ can be promptly understood if we notice that, for positive current polarity, the profile of the self-field produced by the AC current matches the profile of the external inhomogeneous field. As a result, a large amount of vortices is present in the superconductor, meaning that a small current amplitude is sufficient to start the flux-flow. In contrast, for negative current polarity, the profiles of the inhomogeneous field and the AC current self-field are opposites. This means that there are no vortices in the system, unless the current amplitude is large enough to overcome the Bean-Livingston barrier. This is depicted at the second orange arrow in Fig. 9 which occurs at a large current amplitude, very close to $I_{c}^{-}$. This discrepancy is the origin of the large value of $\epsilon_{FFD}$. It is worth mentioning that $I_{cv}$ is not well-defined in cases where the $VI$ characteristics is dominated by thermally activated processes. In such cases $\epsilon_{NSD}$ might be a more reliable measure of the diode efficiency.
@@ III.5 Optimizing diode efficiencies: the role of heat removal and sweep rate
So far, when solving the heat diffusion equation, we have set $\eta=2\times 10^{-4}$, which mimics a substrate of moderate heat removal efficiency. Once we have seen that the destruction of the superconducting state is governed by the emergence and expansion of hot spots through the system, it is natural to assume that the efficiency of our diode system depends on the heat removal coefficient $\eta$. Fig. 11 shows both $\epsilon_{NSD}$ and $\epsilon_{FFD}$ as a function of $\eta$, for values raging from $2\times 10^{-5}$ (weak heat removal) to $2\times 10^{-3}$ (strong heat removal).
@@
Figure 11: Efficiencies of the superconducting, $\epsilon_{\text{NSD}}$ (blue circles), and vortex, $\epsilon_{\text{FFD}}$ (red circles), diode effects as a function of the heat transfer coefficient $\eta$. Here, we set $\tau=10^{6}t_{GL}$.
@@
The vortex diode efficiency presents a weak dependence on $\eta$, which is a result of the low heat dissipation produced by vortices when driven by currents just above $I_{cv}^{+}$ (see the low flux-flow resistance regime shown in Fig. 12). However, as the current increases above $I_{cv}^{+}$, the rate of v-av collisions at the equatorial line of the superconducting film increases, leading to the formation of hot spots as discussed in the previous subsection. For high $\eta$, weak hot spots are quickly removed by the substrate so that a current much larger than $I_{cv}^{+}$ is necessary to drive the system into the normal state. In contrast, for small $\eta$, even weak hot spots, created in the early stages of the flux-flow regime, can grow and drive the film into the normal state. This can considerably reduce the value of $I_{c}^{+}$, bringing it closer to $I_{cv}^{+}$. On the other hand, during the negative current branch, the reduction of $I_{c}^{-}$ induced by weak heat removal is much less pronounced because $I_{c}^{-}$ is already very close to $I_{cv}^{-}$, even in the strong heat removal regime. Combined, these results lead to a pronounced enhancement of $\epsilon_{NSD}$ from 0.067 to 0.317 as $\eta$ is reduced from $2\times 10^{-3}$ to $2\times 10^{-5}$. Therefore, we can conclude that, to increase its efficiency a superconducting diode system should have the weakest possible heat removal scenario.
@@
It is worth mentioning that our 2D simulations are limited to the assumption that the thermal healing length is smaller than the combined thicknesses of film and substrate. In actual experiments, this condition is not generally fulfilled and typical values of $\eta$ can be much smaller than those used in our simulations [46]. In those cases, hot spots can nucleate as soon as flux-flow commences, thus aligning $I_{c}^{+}$ to $I_{cv}^{+}$ and thereby $\epsilon_{NSD}$ to $\epsilon_{FFD}$. Bearing this in mind, we speculate that the sharp superconductor-to-normal transitions observed in Ref. [16] are probably triggered by hot spots induced by the fast penetration of vortices or vortex-antivortex annihilation events in a scenario of very small $\eta$.
@@
Another important factor that influences the efficiency of the diode device is the total sweep time $\tau$, which determines the sweep rate along the current cycle. Fig. 12 shows both $\epsilon_{NSD}$ and $\epsilon_{FFD}$ as a function of $\tau$, with $\eta=2\times 10^{-5}$ fixed. As can be seen, both efficiencies present similar behavior, rapidly increasing for small values of $\tau$ and then approaching a constant value for large values. By increasing $\tau$, we observe a decrease in both $I_{c}^{\pm}$ and $I_{cv}^{\pm}$. The large efficiency occurs because $I_{c}^{+}$ ($I_{cv}^{+}$) decreases in a larger rate than $I_{c}^{-}$ ($I_{cv}^{-}$). In this case, a faster increase rate of the current, i.e., a lower value of $\tau$, leads to larger $I_{c}$ and $I_{cv}$, once there is not enough time for the system to reach the stable state for each region of current values.
@@
Figure 12: Efficiencies of the superconducting, $\epsilon_{\text{NSD}}$ (blue circles), and vortex, $\epsilon_{\text{FFD}}$ (red circles), diode effects as a function of total sweep time $\tau$. Here, $\eta=2\times 10^{-5}$.
@@
III.6 Optimizing diode efficiencies: the role of the inhomogeneous field amplitude and shape